# Geometric Transformations on Null Curves in the Anti-de Sitter $3$-Space

### v1

@@ 4.1. Bäcklund Transformation for the KdV
The Wahlquist-Estabrook equation (4.3) is an overdetermined system whose compatibility equation is the KdV equation (4.1). As a consequence, this system is locally solvable, in the sense that for every $(s_{o},t_{o})\in J\times I\subseteq\mathbb{R}^{2}$ and every constant $c\in\mathbb{R}$ there locally exists a unique solution of (4.3) with initial condition $f(s_{o},t_{o})=c$.
We will next describe a procedure to construct this solution. This method will employ the extended frames of solutions of the KdV equation ([11]).
For a function $\kappa:J\times I\subseteq\mathbb{R}^{2}\longrightarrow\mathbb{R}$ and a constant $\lambda\in\mathbb{R}$, define the $\mathfrak{sl}(2,\mathbb{R})$-valued $1$-form
(4.4) $\Gamma_{\lambda}=\mathcal{K}_{\lambda}\,ds+\mathcal{P}_{\lambda}\,dt\,,$
@@
(4.5) $\mathcal{K}_{\lambda}=\begin{pmatrix}0&\kappa+\lambda\\ 1&0\end{pmatrix},\quad\quad\mathcal{P}_{\lambda}=\begin{pmatrix}-\partial_{s}\kappa&-\partial_{s}^{2}\kappa+2\kappa^{2}-2\lambda\kappa-4\lambda^{2}\\ 2\kappa-4\lambda&\partial_{s}\kappa\end{pmatrix}.$
The $1$-form $\Gamma_{\lambda}$ satisfies the Maurer-Cartan compatibility equation if and only if the function $\kappa(s,t)$ satisfies the KdV equation (4.1). Consequently, as shown in [15], for a given solution $\kappa$ of (4.1) and every $\lambda\in\mathbb{R}$ there exists a map $E_{\lambda}:J\times I\subseteq\mathbb{R}^{2}\longrightarrow{\rm SL}(2,\mathbb{R})$ such that
(4.6) $dE_{\lambda}=E_{\lambda}\,\Gamma_{\lambda}\,,\quad\quad\quad E_{\lambda}(0,0)={\rm Id}\,.$
The maps $E_{\lambda}$ depend in a real analytic fashion on $\lambda\in\mathbb{R}$. The map $E_{\lambda}$ is called an extended frame of $\kappa$ with spectral parameter $\lambda\in\mathbb{R}$ ([11]) [^4] Observe that in the paper [11], the spectral parameter $\lambda$ is a complex number, while here we are restricting it to real values..
@@
(4.7) $\begin{pmatrix}x\\ y\end{pmatrix}=E_{\lambda}^{-1}E_{\lambda}(s_{o},t_{o})\begin{pmatrix}-c\\ 1\end{pmatrix}\,.$
Then, the function $f=-x/y$ is a solution [^5] The function $f=-x/y$ is a local solution of (4.3). Indeed, this solution is singular at the zero locus of the function $y$. of the Wahlquist-Estabrook equation (4.3) with initial condition $f(s_{o},t_{o})=c$ ([11]).
@@
If $f(s,t)$ is a solution of the Wahlquist-Estabrook equation (4.3), then for every $t\in I$ fixed the function $f_{t}(s)$ is a solution of the Riccati equation (3.14).
4.2. Geometric Realization
@@
Consider a smooth one parameter family of null curves $\gamma:(s,t)\in J\times I\subseteq\mathbb{R}^{2}\longmapsto\gamma(s,t)=\gamma_{t}(s)\in{\rm AdS}$ without inflection points and parameterized by the proper time. In other words, for each $t\in I$, we have a null curve $\gamma_{t}:J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ satisfying the assumptions of Remark 2.2.
@@
(4.8) $\partial_{t}\gamma=2\partial_{s}^{3}\gamma-6\kappa\partial_{s}\gamma\,.$
@@
Theorem 4.2 (Theorem 4.2 of [8]).
Let $\gamma:J\times I\subseteq\mathbb{R}^{2}\longrightarrow{\rm AdS}$ be a solution of the LIEN flow (4.8). Then, the bending $\kappa(s,t)$ of $\gamma(s,t)$ evolves according to the KdV equation (4.1). Conversely, if $\kappa:J\times I\subseteq\mathbb{R}^{2}\longrightarrow\mathbb{R}$ is a smooth solution of the KdV equation (4.1), then
$\gamma=E_{1}E_{-1}^{-1}:J\times I\subseteq\mathbb{R}^{2}\longrightarrow{\rm AdS}\,,$
@@
Let $\gamma:J\times I\subseteq\mathbb{R}^{2}\longrightarrow{\rm AdS}$ be a solution of the LIEN flow (4.8) with bending $\kappa(s,t)$ and denote by $(F_{+},F_{-})$ the spinor frame field along $\gamma(s,t)$. Given a constant $\xi\neq 0$ and a solution $f:J\times I\subseteq\mathbb{R}^{2}\longrightarrow\mathbb{R}$ of the Wahlquist-Estabrook equation (4.3) for $\lambda=\cosh(2\xi)$, define the map $\widetilde{\gamma}:J\times I\subseteq\mathbb{R}^{2}\longrightarrow{\rm AdS}$ by
(4.9) $\widetilde{\gamma}=\pm F_{+}\begin{pmatrix}\tanh(\xi)&-{\rm csch}(\xi){\rm sech}(\xi)f\\ 0&\coth(\xi)\end{pmatrix}F_{-}^{-1}\,.$
Then:
  1. (1)
@@
  2. (2)
    The map $\widetilde{\gamma}:J\times I\subseteq\mathbb{R}^{2}\longrightarrow{\rm AdS}$ is a solution of the LIEN flow (4.8) where its bending $\widetilde{\kappa}$ is the Bäcklund transform of $\kappa$ with spectral parameter $\lambda$ and transforming function $f$.
@@
Suppose that $\gamma$ is a solution of the LIEN flow (4.8) and that $f$ satisfies the Wahlquist-Estabrook equation (4.3). As noticed in Remark 4.1, for every $t\in I$, the function $f_{t}(s)$ is a solution of the Riccati equation (3.14). It then follows from Theorem 3.8 that, for every $t\in I$ fixed, the null curve $\widetilde{\gamma}_{t}(s)=\widetilde{\gamma}(s,t)$ defined on (4.9) of the statement is a $\mathcal{T}$-transform of the null curve $\gamma_{t}(s)=\gamma(s,t)$ with parameter $\xi$ and transforming function $f_{t}$. In addition, the bending $\widetilde{\kappa}_{t}$ of $\widetilde{\gamma}_{t}$ is
@@
where $\kappa_{t}$ is the bending of $\gamma_{t}$. Clearly, since $\lambda=\cosh(2\xi)$, $\widetilde{\kappa}_{t}$ is the Bäcklund transform of $\kappa_{t}$ with spectral parameter $\lambda$ and transforming function $f_{t}$. As mentioned above, the Bäcklund transform $\widetilde{\kappa}_{t}$ is also a solution of the KdV equation (4.1). Hence, we deduce from Theorem 4.2 (see also Remark 4.3) and the expressions of $\widetilde{F}_{+}$ and $\widetilde{F}_{-}$ given in (3.6)-(3.7), that $\widetilde{\gamma}:J\times I\subseteq\mathbb{R}^{2}\longrightarrow{\rm AdS}$ is a solution of the LIEN flow. ∎
@@
Let $\gamma:J\times I\subseteq\mathbb{R}^{2}\longrightarrow{\rm AdS}$ be a solution of the LIEN flow (4.8) with bending $\kappa(s,t)$ and consider two $\mathcal{T}$-transforms of $\gamma$, namely, $\gamma_{1}=\mathcal{T}_{\xi_{1},f_{1}}(\gamma)$ and $\gamma_{2}=\mathcal{T}_{\xi_{2},f_{2}}(\gamma)$ with parameters $\xi_{1}\neq\xi_{2}$ and transforming functions $f_{1}\neq f_{2}$, respectively. Then, the function [^6] Observe that the functions $f_{2\to 1}$, $f_{1\to 2}$, $\kappa_{2\to 1}$ and $\kappa_{1\to 2}$ of the statement of Theorem 4.6 are functions of two variables $(s,t)\in J\times I\subseteq\mathbb{R}^{2}$, while the analogue ones of Theorem 3.10 are functions of just one variable $s\in J\subseteq\mathbb{R}$. For simplicity, we avoid explicitly describing this in the statement. $f_{2\to 1}$ (defined as in Theorem 3.10) satisfies the Wahlquist-Estabrook equation (4.3) for $\kappa_{1}$ and $\lambda_{2}=\cosh(2\xi_{2})$, while $f_{1\to 2}$ (also defined as in Theorem 3.10) satisfies (4.3) for $\kappa_{2}$ and $\lambda_{1}=\cosh(2\xi_{1})$. Moreover,
$\mathcal{T}_{\xi_{2},f_{2\to 1}}\left(\mathcal{T}_{\xi_{1},f_{1}}(\gamma)\right)=\mathcal{T}_{\xi_{1},f_{1\to 2}}\left(\mathcal{T}_{\xi_{2},f_{2}}(\gamma)\right),$
and the bending of $\mathcal{T}_{\xi_{2},f_{2\to 1}}\left(\mathcal{T}_{\xi_{1},f_{1}}(\gamma)\right)$ is the function $\kappa_{2\to 1}=\kappa_{1\to 2}$ (defined as in Theorem 3.10). In particular, the bending $\kappa_{2\to 1}$ is a solution of the KdV equation (4.1).
@@
The proof is analogous to that of Theorem 3.10. The only remaining part to prove is that $f_{2\to 1}$ and $f_{1\to 2}$ satisfy the suitable Wahlquist-Estabrook equations (4.3). However, this is a consequence of the classical permutability theorem for the Bäcklund transform of the KdV equation (4.1). ∎
5. Construction Procedure
@@
Consider a constant solution $\kappa$ of the KdV equation (4.1). We first build the extended frames $E_{\lambda}$ of $\kappa$ with spectral parameter $\lambda\in\mathbb{R}$. To this end, we must integrate the $\mathfrak{sl}(2,\mathbb{R})$-valued $1$-form $\Gamma_{\lambda}$ given in (4.4). Since $\kappa$ is constant, it follows from the Maurer-Cartan compatibility equation $d\Gamma_{\lambda}+\Gamma_{\lambda}\wedge\Gamma_{\lambda}=0$ that $\mathcal{K}_{\lambda}$ and $\mathcal{P}_{\lambda}$ commute (see (4.5) for their definition). Therefore, the map $E_{\lambda}(s,t)={\rm Exp}(\mathcal{K}_{\lambda}s+\mathcal{P}_{\lambda}t)$ satisfies (4.6) and so it is an extended frame of $\kappa$ with spectral parameter $\lambda\in\mathbb{R}$. A computation involving the definition of $\mathcal{K}_{\lambda}$ and $\mathcal{P}_{\lambda}$ given in (4.5) shows that
@@
where
$\sigma(s,t)=\sqrt{\kappa(s,t)+\lambda}\left(s+2\left[\kappa(s,t)-2\lambda\right]t\right).$
Here, we are understanding that $\sinh(ix)=i\sin(x)$ and $\cosh(ix)=\cos(x)$.
We will next construct a Bäcklund transform of $\kappa$ with spectral parameter $\lambda\neq 0$ and transforming function $f$. Recall that the transforming function is a solution of the Wahlquist-Estabrook equation (4.3). Hence, we will use the method explained in Subsection 4.1 to construct this solution. If $c\in\mathbb{R}$ is the initial condition, from (4.7) we deduce that the solution $f=-x/y$ of the Wahlquist-Estabrook equation (4.3) is
(5.1) $f(s,t)=\frac{c\sqrt{\kappa+\lambda}\,\cosh\sigma(s,t)+(\kappa+\lambda)\sinh\sigma(s,t)}{\sqrt{\kappa+\lambda}\,\cosh\sigma(s,t)+c\,\sinh\sigma(s,t)}\,.$
@@
(5.2) $\widetilde{E}_{\omega}(s,t)=\frac{1}{\omega-\lambda}R(-f,\lambda,\omega)E_{\lambda}(s,t)R(-f,\lambda,\omega)\,,$
@@
Iterating the process and using again (4.7) to construct a solution of the Wahlquist-Estabrook equation (4.3) as above (but now for $\widetilde{E}_{\omega}$ given in (5.2)), one can compute the transforming function [^7] After the first step, the construction of the extended frames involve only algebraic manipulations, which can be performed with the help of any software of symbolic computations. However, the resulting formulas are very long and, hence, they have been omitted here. $\widetilde{f}$ of $\widetilde{\kappa}$ with spectral parameter $\omega\neq\lambda$ and initial condition $\widetilde{c}$ as well as the corresponding Bäcklund transform $\widehat{\kappa}$ of $\widetilde{\kappa}$, which is a $2$-soliton solution of the KdV equation (4.1).
@@
In the following example we consider the particular case where the original null curve is closed and has constant bending.
@@
Refer to caption   Refer to caption
Figure 5. Left: The transforming function $f_{m,n}$ of $\kappa_{m,n}$ with spectral parameter $\lambda_{p}$ and initial condition $c=0$. Right: The Bäcklund transform $\widetilde{\kappa}_{m,n}$ of $\kappa_{m,n}$ with spectral parameter $\lambda_{p}$ and transforming function $f_{m,n}$. The function $\widetilde{\kappa}_{m,n}$ represents a $1$-soliton solution of the KdV equation (4.1). In both cases $m=4$, $n=1$ and $p=1.4$.
@@
$\omega_{p,r}=p+r+\frac{m^{2}+n^{2}}{m^{2}-n^{2}}\,,\quad\quad\quad r>0\,,$
@@
Refer to caption   Refer to caption
Figure 6. Left: The transforming function $\widetilde{f}_{m,n}$ of $\widetilde{\kappa}_{m,n}$ with spectral parameter $\omega_{p,r}$. Right: The Bäcklund transform $\widehat{\kappa}_{m,n}$ of $\widetilde{\kappa}_{m,n}$ with spectral parameter $\omega_{p,r}$ and transforming function $\widetilde{f}_{m,n}$. The function $\widehat{\kappa}_{m,n}$ represents a $2$-soliton solution of the KdV equation (4.1). In both cases, $m=4$, $n=1$, $p=1.4$ and $r=1$.
@@
Let $\xi_{\lambda}\neq 0$ be a constant such that $\cosh(2\xi_{\lambda})=\lambda_{p}$. Then, the $\mathcal{T}$-transform for $\chi=0$ of $\gamma$ with parameter $\xi_{\lambda}$ and transforming function $f_{m,n}$ is $\mathcal{T}_{\xi_{\lambda},f_{m,n}}(\gamma)$ (its explicit expression is given in (4.9) of Theorem 4.4). Since $\widetilde{\kappa}_{m,n}$ is a traveling wave solution of the KdV equation (4.1), it follows that $\mathcal{T}_{\xi_{\lambda},f_{m,n}}(\gamma)$ is a solution of the LIEN flow (4.8) consisting on the evolution of the initial condition $\gamma_{m,n}$ by rigid motions. Furthermore, for every $t\in I$, the function $s\in\mathbb{R}\longmapsto\widetilde{\kappa}_{m,n}(s,t)-\kappa_{m,n}$ is smooth and rapidly decaying (see the picture on the left of Figure 7). This implies that, for every $t\in I$, the null curve $s\in\mathbb{R}\longmapsto\mathcal{T}_{\xi_{\lambda},f_{m,n}}(\gamma(s,t))$ tends asymptotically to two closed null curves with constant bending $\kappa_{m,n}$ as $s\to\pm\infty$. More precisely, in practical terms, this curve is made up of two disjoint, but congruent, torus knots connected by an arc (see Figure 8).
Refer to caption   Refer to caption
@@
Figure 7. Left: The graph of the rapidly decaying function $s\in\mathbb{R}\longmapsto\widetilde{\kappa}_{m,n}(s,0)-\kappa_{m,n}$. On the left of the yellow point and on the right of the red point $\lvert\widetilde{\kappa}_{m,n}(s,0)-\kappa_{m,n}\rvert\lesssim 2.31\times 10^{-9}$. Right: The graph of the function $s\in\mathbb{R}\longmapsto\widehat{\kappa}_{m,n}(s,0)-\kappa_{m,n}$. On the left of the yellow point and on the right of the red point $\lvert\widehat{\kappa}_{m,n}(s,0)-\kappa_{m,n}\rvert\lesssim 1.73\times 10^{-8}$. These graphs have been computed for the values $m=4$, $n=1$, $p=1.4$ and $r=1$.
Refer to captionRefer to captionRefer to caption
@@
Figure 8. Left: The null curve $s\in\mathbb{R}\longmapsto{\mathcal{T}}_{\xi_{\lambda},f_{m,n}}(\gamma(s,0))$. Center: The two disjoint, but congruent torus knots. Right: The arc connecting the two torus knots. For this example, we have fixed $m=4$, $n=1$ and $p=1.4$.
@@
Finally, we build $\mathcal{T}$-transforms for $\chi=0$ of $\mathcal{T}_{\xi_{\lambda},f_{m,n}}(\gamma)$. For this, we fix $r>0$ and a constant $\xi_{\omega}\neq 0$ such that $\cosh(2\xi_{\omega})=\omega_{p,r}$. Using (5.2) we compute the extended frames $\widetilde{E}_{\omega}$ of $\widetilde{\kappa}_{m,n}$ and the transforming function $\widetilde{f}_{m,n}$. Then, the ${\mathcal{T}}$-transform for $\chi=0$ of ${\mathcal{T}}_{\xi_{\lambda},f_{m,n}}(\gamma)$ with parameter $\xi_{\omega}\neq 0$ and transforming function $\widetilde{f}_{m,n}$ is given by (4.9) and denoted by $\mathcal{T}_{\xi_{\omega},\widetilde{f}_{m,n}}\left(\mathcal{T}_{\xi_{\lambda},f_{m,n}}(\gamma)\right)$. Moreover, the bending of $\mathcal{T}_{\xi_{\omega},\widetilde{f}_{m,n}}\left(\mathcal{T}_{\xi_{\lambda},f_{m,n}}(\gamma)\right)$ is the $2$-soliton solution $\widehat{\kappa}_{m,n}$ of the KdV equation (4.1). In this case, the evolution is not by rigid motions. However, as in the previous case, the functions $s\in\mathbb{R}\longmapsto\widehat{\kappa}_{m,n}(s,t)-\kappa_{m,n}$ are rapidly decaying (see the picture on the right of Figure 7). Thus, the geometrical structure of the evolving curves is similar to the previous case (see Figure 9).

### v2

@@ 4.1 Bäcklund transformation for the KdV
The Wahlquist–Estabrook equation (4.3) is an overdetermined system whose compatibility equation is the KdV equation (4.1). As a consequence, this system is locally solvable, in the sense that for every $(s_{o},t_{o})\in J\times I\subseteq\mathbb{R}^{2}$ and every constant $c\in\mathbb{R}$ there locally exists a unique solution of (4.3) with initial condition $f(s_{o},t_{o})=c$.
We will next describe a procedure to construct this solution. This method will employ the extended frames of solutions of the KdV equation [11].
For a function $\kappa\colon J\times I\subseteq\mathbb{R}^{2}\longrightarrow\mathbb{R}$ and a constant $\lambda\in\mathbb{R}$, define the $\mathfrak{sl}(2,\mathbb{R})$-valued $1$-form
$\Gamma_{\lambda}=\mathcal{K}_{\lambda}\,{\rm d}s+\mathcal{P}_{\lambda}\,{\rm d}t,$ (4.4)
@@
$\mathcal{K}_{\lambda}=\begin{pmatrix}0&\kappa+\lambda\\ 1&0\end{pmatrix},\qquad\mathcal{P}_{\lambda}=\begin{pmatrix}-\partial_{s}\kappa&-\partial_{s}^{2}\kappa+2\kappa^{2}-2\lambda\kappa-4\lambda^{2}\\ 2\kappa-4\lambda&\partial_{s}\kappa\end{pmatrix}.$ (4.5)
The $1$-form $\Gamma_{\lambda}$ satisfies the Maurer–Cartan compatibility equation if and only if the function $\kappa(s,t)$ satisfies the KdV equation (4.1). Consequently, as shown in [15], for a given solution $\kappa$ of (4.1) and every $\lambda\in\mathbb{R}$ there exists a map $E_{\lambda}\colon J\times I\subseteq\mathbb{R}^{2}\longrightarrow{\rm SL}(2,\mathbb{R})$ such that
${\rm d}E_{\lambda}=E_{\lambda}\,\Gamma_{\lambda},\qquad E_{\lambda}(0,0)={\rm Id}.$ (4.6)
The maps $E_{\lambda}$ depend in a real analytic fashion on $\lambda\in\mathbb{R}$. The map $E_{\lambda}$ is called an extended frame of $\kappa$ with spectral parameter $\lambda\in\mathbb{R}$ [11]. [^4] Observe that in the paper [11], the spectral parameter $\lambda$ is a complex number, while here we are restricting it to real values.
@@
$\begin{pmatrix}x\\ y\end{pmatrix}=E_{\lambda}^{-1}E_{\lambda}(s_{o},t_{o})\begin{pmatrix}-c\\ 1\end{pmatrix}.$ (4.7)
Then, the function $f=-x/y$ is a solution [^5] The function $f=-x/y$ is a local solution of (4.3). Indeed, this solution is singular at the zero locus of the function $y$. of the Wahlquist–Estabrook equation (4.3) with initial condition $f(s_{o},t_{o})=c$ [11].
@@
If $f(s,t)$ is a solution of the Wahlquist–Estabrook equation (4.3), then for every $t\in I$ fixed the function $f_{t}(s)$ is a solution of the Riccati equation (3.11).
4.2 Geometric realization
@@
Consider a smooth one parameter family of null curves $\gamma\colon(s,t)\in J\times I\subseteq\mathbb{R}^{2}\longmapsto\gamma(s,t)=\gamma_{t}(s)\in{\rm AdS}$ without inflection points and parameterized by the proper time. In other words, for each $t\in I$, we have a null curve $\gamma_{t}:J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ satisfying the assumptions of Remark 2.2.
@@
$\partial_{t}\gamma=2\partial_{s}^{3}\gamma-6\kappa\partial_{s}\gamma.$ (4.8)
@@
Theorem 4.2 ([8, Theorem 4.2]).
Let $\gamma\colon J\times I\subseteq\mathbb{R}^{2}\longrightarrow{\rm AdS}$ be a solution of the LIEN flow (4.8). Then, the bending $\kappa(s,t)$ of $\gamma(s,t)$ evolves according to the KdV equation (4.1).
Conversely, if $\kappa\colon J\times I\subseteq\mathbb{R}^{2}\longrightarrow\mathbb{R}$ is a smooth solution of the KdV equation (4.1), then
$\gamma=E_{1}E_{-1}^{-1}\colon\ J\times I\subseteq\mathbb{R}^{2}\longrightarrow{\rm AdS}$
@@
Let $\gamma\colon J\times I\subseteq\mathbb{R}^{2}\longrightarrow{\rm AdS}$ be a solution of the LIEN flow (4.8) with bending $\kappa(s,t)$ and denote by $(F_{+},F_{-})$ the spinor frame field along $\gamma(s,t)$. Given a constant $\xi\neq 0$ and a solution $f\colon J\times I\subseteq\mathbb{R}^{2}\longrightarrow\mathbb{R}$ of the Wahlquist–Estabrook equation (4.3) for $\lambda=\cosh(2\xi)$, define the map $\widetilde{\gamma}\colon J\times I\subseteq\mathbb{R}^{2}\longrightarrow{\rm AdS}$ by
$\widetilde{\gamma}=\pm F_{+}\begin{pmatrix}\tanh(\xi)&-{\rm csch}(\xi){\rm sech}(\xi)f\\ 0&\coth(\xi)\end{pmatrix}F_{-}^{-1}.$ (4.9)
Then,
  1. $1$
@@
  2. $2$
    The map $\widetilde{\gamma}\colon J\times I\subseteq\mathbb{R}^{2}\longrightarrow{\rm AdS}$ is a solution of the LIEN flow (4.8) where its bending $\widetilde{\kappa}$ is the Bäcklund transform of $\kappa$ with spectral parameter $\lambda$ and transforming function $f$.
@@ 2. $2$
Suppose that $\gamma$ is a solution of the LIEN flow (4.8) and that $f$ satisfies the Wahlquist–Estabrook equation (4.3). As noticed in Remark 4.1, for every $t\in I$, the function $f_{t}(s)$ is a solution of the Riccati equation (3.11).
It then follows from Theorem 3.8 that, for every $t\in I$ fixed, the null curve $\widetilde{\gamma}_{t}(s)=\widetilde{\gamma}(s,t)$ defined on (4.9) of the statement is a $\mathcal{T}$-transform of the null curve $\gamma_{t}(s)=\gamma(s,t)$ with parameter $\xi$ and transforming function $f_{t}$. In addition, the bending $\widetilde{\kappa}_{t}$ of $\widetilde{\gamma}_{t}$ is
@@
where $\kappa_{t}$ is the bending of $\gamma_{t}$. Clearly, since $\lambda=\cosh(2\xi)$, $\widetilde{\kappa}_{t}$ is the Bäcklund transform of $\kappa_{t}$ with spectral parameter $\lambda$ and transforming function $f_{t}$.
As mentioned above, the Bäcklund transform $\widetilde{\kappa}_{t}$ is also a solution of the KdV equation (4.1). Hence, we deduce from Theorem 4.2 (see also Remark 4.3) and the expressions of $\widetilde{F}_{+}$ and $\widetilde{F}_{-}$ given in (3.6) and (3.7), that $\widetilde{\gamma}\colon J\times I\subseteq\mathbb{R}^{2}\longrightarrow{\rm AdS}$ is a solution of the LIEN flow. ∎
@@
[…]
@@
The proof is analogous to that of Theorem 3.10. The only remaining part to prove is that $f_{2\to 1}$ and $f_{1\to 2}$ satisfy the suitable Wahlquist–Estabrook equations (4.3). However, this is a consequence of the classical permutability theorem for the Bäcklund transform of the KdV equation (4.1). ∎
5 Construction procedure
@@
Consider a constant solution $\kappa$ of the KdV equation (4.1). We first build the extended frames $E_{\lambda}$ of $\kappa$ with spectral parameter $\lambda\in\mathbb{R}$. To this end, we must integrate the $\mathfrak{sl}(2,\mathbb{R})$-valued $1$-form $\Gamma_{\lambda}$ given in (4.4). Since $\kappa$ is constant, it follows from the Maurer–Cartan compatibility equation ${\rm d}\Gamma_{\lambda}+\Gamma_{\lambda}\wedge\Gamma_{\lambda}=0$ that $\mathcal{K}_{\lambda}$ and $\mathcal{P}_{\lambda}$ commute (see (4.5) for their definition). Therefore, the map $E_{\lambda}(s,t)={\rm Exp}(\mathcal{K}_{\lambda}s+\mathcal{P}_{\lambda}t)$ satisfies (4.6) and so it is an extended frame of $\kappa$ with spectral parameter $\lambda\in\mathbb{R}$. A computation involving the definition of $\mathcal{K}_{\lambda}$ and $\mathcal{P}_{\lambda}$ given in (4.5) shows that
@@
where $\sigma(s,t)=\sqrt{\kappa+\lambda}(s+2[\kappa-2\lambda]t)$. Here, we are understanding that $\sinh({\rm i}x)={\rm i}\sin(x)$ and $\cosh({\rm i}x)=\cos(x)$.
We will next construct a Bäcklund transform of $\kappa$ with spectral parameter $\lambda\neq 0$ and transforming function $f$. Recall that the transforming function is a solution of the Wahlquist–Estabrook equation (4.3). Hence, we will use the method explained in Subsection 4.1 to construct this solution. If $c\in\mathbb{R}$ is the initial condition, from (4.7) we deduce that the solution $f=-x/y$ of the Wahlquist–Estabrook equation (4.3) is
$f(s,t)=\frac{c\sqrt{\kappa+\lambda}\,\cosh\sigma(s,t)+(\kappa+\lambda)\sinh\sigma(s,t)}{\sqrt{\kappa+\lambda}\,\cosh\sigma(s,t)+c\,\sinh\sigma(s,t)}.$ (5.1)
@@
$\widetilde{E}_{\omega}(s,t)=\frac{1}{\omega-\lambda}R(-f,\lambda,\omega)E_{\lambda}(s,t)R(-f,\lambda,\omega),$ (5.2)
@@
Iterating the process and using again (4.7) to construct a solution of the Wahlquist–Estabrook equation (4.3) as above (but now for $\widetilde{E}_{\omega}$ given in (5.2)), one can compute the transforming function [^7] After the first step, the construction of the extended frames involve only algebraic manipulations, which can be performed with the help of any software of symbolic computations. However, the resulting formulas are very long and, hence, they have been omitted here. $\widetilde{f}$ of $\widetilde{\kappa}$ with spectral parameter $\omega\neq\lambda$ and initial condition $\widetilde{c}$ as well as the corresponding Bäcklund transform $\widehat{\kappa}$ of $\widetilde{\kappa}$, which is a $2$-soliton solution of the KdV equation (4.1).
@@
In the following example, we consider the particular case where the original null curve is closed and has constant bending.
@@
Figure 5: Left: The transforming function $f_{m,n}$ of $\kappa_{m,n}$ with spectral parameter $\lambda_{p}$ and initial condition ${c=0}$. Right: The Bäcklund transform $\widetilde{\kappa}_{m,n}$ of $\kappa_{m,n}$ with spectral parameter $\lambda_{p}$ and transforming function $f_{m,n}$. The function $\widetilde{\kappa}_{m,n}$ represents a $1$-soliton solution of the KdV equation (4.1). In both cases, $m=4$, $n=1$ and $p=1.4$.
@@
$\omega_{p,r}=p+r+\frac{m^{2}+n^{2}}{m^{2}-n^{2}},\qquad r>0,$
@@
Figure 6: Left: The transforming function $\widetilde{f}_{m,n}$ of $\widetilde{\kappa}_{m,n}$ with spectral parameter $\omega_{p,r}$. Right: The Bäcklund transform $\widehat{\kappa}_{m,n}$ of $\widetilde{\kappa}_{m,n}$ with spectral parameter $\omega_{p,r}$ and transforming function $\widetilde{f}_{m,n}$. The function $\widehat{\kappa}_{m,n}$ represents a $2$-soliton solution of the KdV equation (4.1). In both cases, $m=4$, $n=1$, $p=1.4$ and $r=1$.
@@
Let $\xi_{\lambda}\neq 0$ be a constant such that $\cosh(2\xi_{\lambda})=\lambda_{p}$. Then, the $\mathcal{T}$-transform for $\chi=0$ of $\gamma$ with parameter $\xi_{\lambda}$ and transforming function $f_{m,n}$ is $\mathcal{T}_{\xi_{\lambda},f_{m,n}}(\gamma)$ (its explicit expression is given in (4.9) of Theorem 4.4). Since $\widetilde{\kappa}_{m,n}$ is a traveling wave solution of the KdV equation (4.1), it follows that $\mathcal{T}_{\xi_{\lambda},f_{m,n}}(\gamma)$ is a solution of the LIEN flow (4.8) consisting of the evolution of the initial condition $\gamma_{m,n}$ by rigid motions. Furthermore, for every $t\in I$, the function ${s\in\mathbb{R}\longmapsto\widetilde{\kappa}_{m,n}(s,t)-\kappa_{m,n}}$ is smooth and rapidly decaying (see the picture on the left of Figure 7). This implies that, for every $t\in I$, the null curve $s\in\mathbb{R}\longmapsto\mathcal{T}_{\xi_{\lambda},f_{m,n}}(\gamma(s,t))$ tends asymptotically to two closed null curves with constant bending $\kappa_{m,n}$ as $s\to\pm\infty$. More precisely, in practical terms, this curve is made up of two disjoint, but congruent, torus knots connected by an arc (see Figure 8).
@@
Figure 7: Left: The graph of the rapidly decaying function $s\in\mathbb{R}\longmapsto\widetilde{\kappa}_{m,n}(s,0)-\kappa_{m,n}$. On the left of the yellow point and on the right of the red point $\lvert\widetilde{\kappa}_{m,n}(s,0)-\kappa_{m,n}\rvert\lesssim 2.31\times 10^{-9}$. Right: The graph of the function $s\in\mathbb{R}\longmapsto\widehat{\kappa}_{m,n}(s,0)-\kappa_{m,n}$. On the left of the yellow point and on the right of the red point $\lvert\widehat{\kappa}_{m,n}(s,0)-\kappa_{m,n}\rvert\lesssim 1.73\times 10^{-8}$. These graphs have been computed for the values $m=4$, $n=1$, $p=1.4$ and $r=1$.
@@
Figure 8: Left: The null curve $s\in\mathbb{R}\longmapsto{\mathcal{T}}_{\xi_{\lambda},f_{m,n}}(\gamma(s,0))$. Center: The two disjoint, but congruent torus knots. Right: The arc connecting the two torus knots. For this example, we have fixed $m=4$, $n=1$ and $p=1.4$.
@@
Finally, we build $\mathcal{T}$-transforms for $\chi=0$ of $\mathcal{T}_{\xi_{\lambda},f_{m,n}}(\gamma)$. For this, we fix $r>0$ and a constant ${\xi_{\omega}\neq 0}$ such that $\cosh(2\xi_{\omega})=\omega_{p,r}$. Using (5.2), we compute the extended frames $\widetilde{E}_{\omega}$ of $\widetilde{\kappa}_{m,n}$ and the transforming function $\widetilde{f}_{m,n}$. Then, the ${\mathcal{T}}$-transform for $\chi=0$ of ${\mathcal{T}}_{\xi_{\lambda},f_{m,n}}(\gamma)$ with parameter $\xi_{\omega}\neq 0$ and transforming function $\widetilde{f}_{m,n}$ is given by (4.9) and denoted by $\mathcal{T}_{\xi_{\omega},\widetilde{f}_{m,n}}\bigl{(}\mathcal{T}_{\xi_{\lambda},f_{m,n}}(\gamma)\bigr{)}$. Moreover, the bending of $\mathcal{T}_{\xi_{\omega},\widetilde{f}_{m,n}}\bigl{(}\mathcal{T}_{\xi_{\lambda},f_{m,n}}(\gamma)\bigr{)}$ is the $2$-soliton solution $\widehat{\kappa}_{m,n}$ of the KdV equation (4.1). In this case, the evolution is not by rigid motions. However, as in the previous case, the functions $s\in\mathbb{R}\longmapsto\widehat{\kappa}_{m,n}(s,t)-\kappa_{m,n}$ are rapidly decaying (see the picture on the right of Figure 7). Thus, the geometrical structure of the evolving curves is similar to the previous case (see Figure 9).